Reports 1-1 of 1 Clear search Modify search
Virgo Runs (O3)
berni - 23:00 Monday 01 April 2019 (45502) Print this report
Operator report - afternoon shift
ITF found in commissioning mode for the last minute activities.

At 15:00 UTC ITF in Science mode for the start of O3 run.

From 20:30 UTC to 20:40 UTC we had a series of glitches in the horizon, see attached plot.

Environment
- Rosario worked at BcMEteo, BcLight server; the servers are running but data not acquired yet.
- earthquake at 18:37 UTC with magnitude 5.5; see attached picture.

Suspensions
Yellow flag for NI_IP - Sa_NI_F0_COIL_H2_50Hz (coil close to saturation). Experts are aware of the situation and it will be fixed tomorrow morning.

Guard tours
no info received from the guardian.
Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×